OpenStudy (anonymous):

You decide to see if you can ride the elevator to street level (Floor 0) without pushing any buttons. The elevator takes you up four floors, down 6 floors, up 1 floor, down 8 floors, down 3 floors, up 1 floor, and then down 6 floors to street level. Write an equation to find your starting floor.

OpenStudy (anonymous):

Well, you start from street level, then reverse up and down. Street level = 0 down = + up = - Answer = 0 + 6 - 1 + 3 + 8 - 1 + 6 - 4 = 17 You started to ride the lift on 17th floor

OpenStudy (anonymous):

Awesome, thanks very much!

OpenStudy (anonymous):

Can anyone help me by putting this into an algebraic expression?

OpenStudy (anonymous):

Where x is the floor you start from, x = 0 + 6 - 1 + 3 + 8 - 1 + 6 - 4 x = 17

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

This is how the problem expressed in algebra equation: x + 4 - 6 + 1 - 8 - 3 + 1 - 6 = 0 Try to solve that, answer should be the same.

OpenStudy (anonymous):

Yup, worked perfectly! Thanks for explaining :o)

OpenStudy (anonymous):

Just posted a big one if you have time to see if you can assist.